Chapter 05 Homework omework 6 Northwood Company manufactures basketballs. The company has a small plant that relies heavily on direct labor workers. Thus, variable expenses are high, totaling $1500 per ball, of which 60% is direct ball that sells for $25. At present, the bal is manufactured in a t labor cost Last year,the company sold 44,000 of these balls, with the following results: Sales (44,000 balla) Variable expenses Contribution margin Pixed expenses 1.100,000 Net operating ineome Required: 2. Due to an increase in labor rates, the company estimates that next years variable expenses will nc change takes place and the selling price per ball remainsc point in balls? leverage at last years sales level. rease by $3.00 per balil If this nces 3efer to the deta in (2) above. If the expected change in variable expenses takes place, how many balls will have to be sold next company must raise the selling price of its basketbals. f year to earn the same net operating income, $123,000, as last yeer? again to the data in (2) above. The president feels that the t Company wants to maintain the same CM ratio as last year (as computed in requirement la), w hat selling price per bal must it charge next year to cover the increased labor costs? would slash variable expenses per bal by 4000%, but it would would be the companys new CM ratio and new bresk-even point in bals? r to the original data. The company is discussing the construction of a new, automated cause fixed expenses per year to double. if the new plant is built, what 6. Refer to the diata in (5) above. n. If the new plant is built, how many balls will have to be sold next year to earn the same net operating income, $123,000, as last year? b. Assume year), Prepare a contribution format income statement and Compute the degree of operating leverage the new plant is built and that next year the company manufactures and sells 44,000 balls (the same number as sold last Complete this question by entering your answers in the tabs below. Prev 3014 İİİ Next> 2 3 4
media%2Fec6%2Fec6ac276-75e7-4120-969a-a7
media%2F6b0%2F6b089711-19e3-407c-9496-c9
media%2Ff14%2Ff1460523-ecb1-49d4-97dd-8c
media%2F628%2F628e8712-528d-417d-b905-26
media%2F4da%2F4da169cc-19f9-4411-91ec-76
media%2F1c9%2F1c959e1d-d84f-493b-9b31-1f
media%2F22f%2F22f77e90-d95d-4cd0-b8d8-b7
Answer #1

1

Contribution Margin ratio

$

a

contribution

440000

b

sales

1100000

c

(a/b)

40%

Break even point:

a

Fixed expenses

317000

b

Contribution Margin per unit

10

c

(a/b)

31700 units

Operating leverage:

a

Contribution Margin

440000

b

Net operating income

123000

c

(a/b)

3.58

2

New contribution ratio will be

sales

25

100%

less: Variable cost [15+3]

18

72%

contribution Margin

7

28%

Break even point

[317000/7]

45285.71

3

sales

25

100%

less: Variable cost [15+3]

18

72%

contribution Margin

7

28%

contribution(profit+ fixed expense)

(123000+317000)

440000

sales-variable cost = fixed cost + desired profit

25X-18X = 440000

x = 440000/7

62857.14

4

CM ratio = [(Sales price *units) - New variable units]/(sales price*units)

0.4 = [44000x - (44000*18)]/44000x

17600x = [44000X-792000]

44000X+17600X = 792000

61600X = 792000

X = 792000/61600

X = $12.85